OpenStudy (nicoleced):

i have another problem i need detailed explanation V =1/3 Ah

OpenStudy (nicoleced):

solve for h

OpenStudy (anonymous):

Ok Well V Stands for Viva, and the rest stands for Columbia. So it would equal Viva Columbia!

OpenStudy (anonymous):

Sorry idk.

OpenStudy (nicoleced):

ITS COLOMBIA for the billionth time! ugh 😒

OpenStudy (anonymous):

Jeezus.

OpenStudy (nicoleced):

I need help with my math problem. somebody please? :)

OpenStudy (anonymous):

V = (1/3) A h 3V = A h h = 3V / A

OpenStudy (anonymous):

can you help me

OpenStudy (nicoleced):

i think... 1/3 = 3 so 3v= ah so then you divide and h= 3v/a

OpenStudy (nicoleced):

right?

OpenStudy (anonymous):

you multiply three to to v and then divide a from h

OpenStudy (nicoleced):

got it.

OpenStudy (nicoleced):

what do you need help with ?

OpenStudy (nicoleced):

thank you.